Aina and Joan invest in each account.Aina invests Php 275,000 at 10% for 15 years and joan invests Php 275,000 at 12% for 12 years.Both investments are compounded annually.Which one will yield more?


1
Expert's answer
2021-05-07T09:49:49-0400

Compound Interest:

"A = P(1 + \frac{r}{n})^{nt}"

P: the principal, amount invested

A: the new balance

t: the time

r: the rate, (in decimal form)

n: the number of times it is compounded.

Aina:

P = 275000

t = 15

r = 0.10

compounded annually, n = 1

"A = 275000(1 + \frac{0.10}{1})^{1 \times 15} = 1148743.24"

Joan:

P = 275000

t = 12

r = 0.12

compounded annually, n = 1

"A = 275000(1 + \frac{0.12}{1})^{1 \times 12} = 1071393.39"

Aina will yield more.